OpenStudy (anonymous):

Cleo and Clare are looking from their balcony to a swimming pool below that is located 15 m horizontally from the bottom of their building. They estimate the balcony is 45 m high and wonder how fast they would have to jump horizontally to succeed in reaching the pool. What calculations would you show to help them determine the answer? Evaluate the practicality of their being able to succeed at jumping into the pool.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

In three scenarios a truck is towing a boat on a trailer. Using 3 – 4 complete sentences indicate which scenario(s) (if any) is the force of the truck on the trailer greater than the force of the trailer on the truck: Scenario A: masstruck = 1000 kg, massboat and trailer = 500 kg; velocity = 20 m/s Scenario B: masstruck = 800 kg, massboat and trailer = 800 kg; velocity = 20 m/s Scenario C: masstruck = 500 kg, massboat and trailer = 1000 kg; velocity = 15 m/s

OpenStudy (michele_laino):

hello!

OpenStudy (michele_laino):

have you got any ideas?

OpenStudy (anonymous):

Not at all, i got some some earlier but theres a few that I'm lost on.

OpenStudy (michele_laino):

I think that in all of your three cases, both forces have the same magnitude, since we have to apply the third law of Newton

OpenStudy (michele_laino):

furthermore, since the velocity is constant, then we can consider the system truck+boat, an inertial system

OpenStudy (michele_laino):

and, as you know, the laws of Physics are the same into all inertial systems

OpenStudy (anonymous):

so none of them are greater than the other? correct?

OpenStudy (michele_laino):

yes! correct, since the velocity of the system truck+boat, is constant in all of your three cases

OpenStudy (anonymous):

how could i try and explain that in a few sentences?

OpenStudy (michele_laino):

here is my sentence: "since the velocity of the system truck+boat" is constant, as vector, then the same system "truck+boat" is an inertial system. Now within an inertial system, all law of Physics, have the same shape, as formula, so we can apply the third law of Newton, and we can conclude that the force applied by truck on boat have the same direction and magnitude of the force applied by the boat on truck. Furthermore those forces have opposite orientation each other"

OpenStudy (anonymous):

what about the first question?

OpenStudy (michele_laino):

we have to consider these two equations: \[\left\{ \begin{gathered} h = \frac{1}{2}g{t^2} \hfill \\ \hfill \\ d = vt \hfill \\ \end{gathered} \right.\] where d=15 meters and h= 45 meters

OpenStudy (michele_laino):

now, I solve the second equation for t, and I get: \[t = \frac{d}{v}\]

OpenStudy (michele_laino):

Now I substitute into the first equation, and I get: \[45 = \frac{1}{2}g{\left( {\frac{{15}}{v}} \right)^2}\] please solve that equation for velocity v, here g=9.81 m/sec^2

OpenStudy (anonymous):

i dont understand you math I'm confused

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

the subsequent drawing represents the situation of your exercise: